OpenStudy (anonymous):

Could someone please help? I do not think I'm doing this right :( or maybe I am missing steps. This question is multiple choice but I am not sure how to get it to look like the answer choices. Solve. 2x^2-8x-12=0

OpenStudy (aravindg):

ok so here you have to solve the quadratic

OpenStudy (anonymous):

Like this? 2x^2-8x-12=0 +8x+12x +8x+12x 2x^2+8x+12x=8x+12x

OpenStudy (aravindg):

use the quadratic formula to get value for x ie for a quadratic \(ax^2+bx+c\) \[x=\dfrac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (aravindg):

* \[ax^2+bx+c=0\]

OpenStudy (anonymous):

x^2=4x+6?

OpenStudy (nubeer):

\[x= \frac{ -b \pm \sqrt{? b^{2} -4ac}}{ ? }\] in your question 2x^2-8x-12=0 a=2, b = -8, c =-12 .... just plug in values u will get 2 answers.

OpenStudy (nubeer):

sorry seems like i missed 2a in denominator.. sorry

OpenStudy (aravindg):

lols :P

OpenStudy (aravindg):

well after you divide by 2 you will get \[x^2-4x-6=0\]

OpenStudy (aravindg):

ok use quadratic formula now

OpenStudy (anonymous):

x=-8±√(8^2-4(2)(12)/4)

OpenStudy (aravindg):

a=1 , b=-4 ,c=-6

OpenStudy (aravindg):

look at the values above ^

OpenStudy (aravindg):

substitute it in quadratic equation

OpenStudy (aravindg):

@P.nut1996 do it !

OpenStudy (anonymous):

My answer choices are x = 1 ± 2 x = 2 ± x = 2 ± x = 2 ± 2

OpenStudy (aravindg):

what did u get after applying quadratic formula?